From GlueXWiki
Jump to: navigation, search

February 27, 2014 FDC meeting


  1. Installation status [1]
  2. Engineering (Bill, Dave)
    1. Cathode scan
    2. LV fast shutdown
  3. Electronics problematic channels (Fernando, Chris, Nick)
  4. Mini-DAQ status FDC E-log (Beni)
  5. Other


Participants: Eugene, Bill, Dave, Nick, Chris, Simon, Beni, and Lubomir.


- A cable holder frame has been installed at the top on face of the magnet (picture attached). It's too far from the mesh and the cables will sag (Nick), but better than before. We may need such frames also on the sides. New cable tray will be installed on the right side of the magnet to avoid crossing with the BCAL water pipes. Tom is working on the cable cradle that will slide together with the FDC on bearings (Bill). Bill is manufacturing the FDC-to-mesh connections made of metal and plastic to insulate the two. We also have to cut the mesh loop. Need to have also caps for the Fluorinert tubes to be used when the cooling system is disconnected from the FDC.

- Yesterday Lubomir replaced the mass flow controller (6000 sccpm range) of the temporary gas system with a re-calibrated one of 500 sccpm. It turned out that to have the same flow on the rotameter we had to set the flow at 200 sccpm or less, to be compared to setting of the previous flow controller of 400sccpm. It turned out also that gas bottle pressure goes down slower than before, but we need more days to figure this out. If this is the case then the old flow controller was leaky, but at the same time it is strange that we are flushing now successfully all the four packages with just 200sccpm total; normally we needed such a flow for one package only.


- Bill returned from Hexagon Metrology, Charlotte, NC, and showed us pictures of the MCC machine and a lot of data: the xlsx file is attached above containing all the raw data. The cathode was scanned optically and automatically with a small step (5mm?) along the strips recording the edges of the strips. Sometimes the machine re-scanned the same point, sometime was obstructed by a defect, or by the tape on the back of the cathode that connects the foils. In any case, this is not a problem since the number of points is huge. Eugene proposed for a better visualization to make a scattered plot of all the (x,y) points. This will help to decide how to analyze the data.

- Dave figured out how to get the status signals from the chiller: it turned out the schematics in the chiller documentation were not up to date. Dave had already everything connected, waiting only for a power supply, so maybe tomorrow we will be able to test the fast shut down. First Hovanes will try it on a different crate. Dave received the flow sensor for the chilled water but not yet the PLS module to be used for the flow monitoring.


- Nick has a spare F1TDCV3 module so we can install it right away.

- Nick and Chris fixed one very bad flat cable connector with mixed wires at the detector side. There are still several 3M connectors to be fixed, but we want to make sure with the mini-DAQ that the problem is there. According to Chris may take many hours to fix such a connector.


- Beni: according to Cody the new and working firmware with the trigger window set by programmable parameters should be ready by the end of the week. Then we need from Bryan Moffit the software to set up the parameters. Eugene: this work is on the critical path now and everybody should know it has high priority. In case the firmware is not ready by the beginning of the next week then we will ask Cody to have these parameters hard coded. If this doesn't work we may consider going back to the Gerard's firmware that is slow but operational, however we don't want to go that back (Eugene).